# IoT Ground Sensing Systems for Early Wildfire Detection: Technologies, Challenges and Opportunities

### v1

@@ Abstract
Wildfire, also widely known as bushfire in Australia and forest fire in Europe, present a significant risk to people, the environment and assets. In recent decades, the surface air temperature in Australia has risen by C due to climate change, leading to more days with high fire risk [1]. As fire weather worsens, fire occurrence increases [2, 3] and the window of opportunity for successful containment and suppression after ignition narrows [4, 5]. Hence, rapid detection of new ignition will become an increasingly important strategy for fire managers as the climate warms.
@@ I-A Impact of Wildfire Across the Globe
In Australia, this recently reached a catastrophic climax in the 2019-2020 “Black Summer” fires across multiple states where 24 million hectares of land were burnt, causing 33 deaths and approximately 450 more injured by smoke inhalation [6]. The 2019-2020 fires burnt an unprecedented area of land [7] and were likely worsened by climate change [8]. The toll on wildlife was immense. The World Wide Fund for Nature records “143 million mammals, 2.46 billion reptiles, 180 million birds and 51 million frogs” had been harmed, killed, or displaced [9] as a result of the Black Summer. For the already fragile ecosystem of animals in Australia, this is a major hit, as many animals, including koalas, wombats and wallabies, are listed as endangered or vulnerable [10].
@@
Worldwide, wildfires are getting worse, with the European Forest Fire Information System estimating that 837,000 hectares of land have been burnt in countries of the European Union alone in 2022, which is the second worst year for wildfires in its history [11]. The United States also experienced record-setting fires in 2020, with more than 50,000 fires burning nearly 3.64 million hectares of land [12]. Although the number of reported wildfires remains stable in the last 30 years, the area burnt in a year caused by wildfires has increased by 1.62 million hectares compared to the area burnt in the 1990s in the United States [13]. The Brazilian National Institute of Space Research has revealed 3.36 million hectares of land burnt in Brazil annually since 2003 [14]. Notably, in January 2017, the Mediterranean region of Chile experienced hundreds of fires burning more than 500,000 hectares, approximately 14 times the average annual area burnt in the region in the last 40 years [15]. In addition to these observed increases in fire activity, predicted warming and drying of the climate across much of the world are likely to increase fire activity and result in long term declines in forest carbon storage, further warming the climate [16]. In light of the increasing threat posed by wildfires, the imperative to develop precise and reliable detection technologies has taken precedence.
@@
I-B Technologies for Early Wildfire Detection
The ability to detect wildfires in their early stages allows effective and efficient fire suppression [17]. Specifically, early detection of initial fire stages, such as smouldering and ground fire, enables the timely deployment of fire suppression tactics to increase the probability of initial attack success [5]. Therefore, it is vital to develop early wildfire detection methods and technologies. Here, we present a set of requirements to gauge the capability of existing technologies for early wildfire detection. The algorithm should be able to identify fires while they are still small enough to be successfully put out. As the fire grows larger, it becomes eventually too difficult to extinguish [5]. Therefore, early detection that can reduce the time it takes to detect or spot small fires is hugely beneficial. Moreover, a wildfire detection system based on environmental monitoring is highly desirable due to its connection to fire hazards. Finally, the scalability of the wildfire detection system is beneficial so that the system can not only be deployed on a large scale but also be deployed in different terrains.
We hereby briefly introduce the existing technologies for wildfire detection and analyse their suitability for early wildfire detection. For nearly a century, watchtowers and public reporting have been used to detect wildfires. Unfortunately, it requires observers to remain on top of the tower during certain hours of the day and specific periods of the year, constantly looking for smoke that implies fire or the effectiveness of public reporting is significantly influenced by the population density of the region. With the advancement of computer vision technologies [18, 19, 20], this task can now be accomplished day and night with the help of arrays of high-resolution vision / thermal cameras that record visual information, which is then processed using different detection algorithms to recognise smoke, flame, or heat from fire. However, there are some challenges associated with this technology for early wildfire detection. Firstly, fires can smoulder before reaching a size that is detectable from a distance, which can prolong their detection time. Secondly, monitoring hilly terrain becomes difficult due to line-of-sight blockage. Consequently, it may be difficult to detect small fires in these areas. Thirdly, this technology lacks the ability to convey environmental changes, such as changes in relative humidity and air pressure, which cause diurnal changes in fuel moisture and associated fire activity [21].
As another detection technology, the deployment of crewed fire-spotting aircraft over lightning strike areas or areas of high fire risk is also used, but is often delayed due to resourcing issues or restrictions on night-flying operations. The use of long-endurance drones is on the rise, primarily due to their availability to operate in rough weather conditions, day and night, helping to overcome some of the limitations associated with traditional fire-spotting aircraft [22, 23, 24]. Unmanned aerial vehicles (UAVs) offer the potential to rapidly confirm suspected ignitions and provide information on the size and behaviour of fires, particularly fires in remote locations. However, this system requires other monitoring technologies to identify suspected ignitions.
For many years, satellite imagery has been a valuable tool for detecting large-scale wildfires [25]. But it also comes with limitations, particularly in terms of temporal or spatial resolutions, making it challenging to detect fires when they are still small. The emerging Low Earth Orbit satellite constellation network offers relatively high spatial resolution capable of detecting small fires [26], however, the temporal frequency (several hours to days) prevents reliable fire detection [27]. Novel algorithms developed specifically for early fire detection using geostationary satellites may offer advances, but have not yet been thoroughly tested [28].
In addition to the technologies mentioned above, a promising complementary approach to detect early wildfires is to use a ground sensing system based on the Internet of Things (IoT). This system provides real-time data from a variety of ground sensors at desired densities for timely and accurate fire detection, due to the proximity of the ground sensors to the fire ignition locations. This IoT-based system can autonomously raise early alarms for wildfire detection, identify its location, and provide situational awareness by reporting the fire movement. The communication delay from data generation at the ground sensors to reception at the central node can be within a few minutes. Moreover, this kind of system can complement other fire detection and extinguishing systems by identifying the area of interest, which then can be investigated for further details and validation of ignitions as well as to initialise responses. The benefits of a ground sensing system would also go beyond wildfire detection and suppression, as they can also be used for environmental monitoring purposes.
Due to its unique benefits, many IoT companies have integrated IoT-based wildfire detection systems into their offerings [29, 30, 31]. However, it is essential not to rely solely on one detection system for early fire detection, but to recognise the benefits of combining different wildfire detection approaches and take a more comprehensive approach. This requires long-term research and development efforts across academia and industry on a large scale. In Australia, a leading effort on this front is the recent establishment of the Bushfire Research Centre of Excellence at the Australian National University (ANU). The Centre aims to develop an integrated and layered solution to detect and extinguish small fires using multiple technologies, some mentioned above [17, 24].
I-C Paper Contributions
There are only a handful of survey papers on this research topic. In [32, 33], the authors provided a summary of recent early fire detection and prevention technologies, while IoT is only a small part of the survey. In [34, 35, 36], comprehensive studies are presented on IoT-based wildfire detection using Machine Learning (ML) and Deep Learning (DL) approaches. However, the challenges faced in the deployment of IoT sensing systems are only briefly mentioned. Therefore, our objective is to explore the potential benefits of modern sensing technologies and their compatibility with IoT sensing systems to detect early-stage wildfires. Specifically, we investigate the use of vision-based and environmental monitoring technologies, combined with anomaly detection algorithms, to detect wildfires in their early stages. Furthermore, the study highlights critical factors to consider when implementing an IoT-based sensing system for reliable and efficient early wildfire detection, such as sensor placement and network connectivity. The objective of this article is to provide a comprehensive understanding and guidance for the implementation of an IoT ground sensing system for early wildfire detection, together with its challenges and possible research extensions.
The contributions of this work are summarised as follows:
An overview of existing detection techniques for vision and environmental sensing data. This article aids in data analytics-based algorithm development.
Identification of operational challenges associated with the deployment of an IoT system in remote areas.
Presenting a framework for developing an IoT ground sensing system for early wildfire detection. This framework is designed to help decision makers determine the best deployment strategies based on the performance metrics given.
I-D Paper Organisation
The structure of this article is outlined in Figure 1. Section II discusses the architecture of IoT systems and IoT camera modules and environmental monitoring sensors that can be used for such an application. It also provides a brief overview of the operating principles of different types of sensors, their characteristics and use case examples. Finally, the performance of the sensors is compared in multiple aspects. In Section III, we provide an overview of the wildfire detection algorithms commonly used in IoT systems. These algorithms can be divided into two categories: vision-based detection and anomaly detection based on environmental sensing data. We further classify these detection methods based on their complexity, such as feature detection, time series pattern detection and ML & DL detection in vision-based and threshold detection, time series detection, ML & DL detection in environmental sensing. Additionally, we discuss energy consumption, detection delay, requirements for deploying detection algorithms and the popularity of different detection algorithms is also provided at the end of the section. Section IV explains the operational challenges associated with the deployment of IoT applications for wildfire detection, including sensor limitations, energy efficiency considerations, network coverage constraints, the accuracy of detection algorithms and the challenges related to the system deployment. Finally, Section V proposes a framework for the deployment of IoT ground sensing systems in the context of early wildfire detection, offering solutions to most of the challenges identified in Section IV. Recognising the inherent trade-offs among some of these challenges, this section also introduces a performance metric that can guide IoT designers in their decision-making processes. Furthermore, it explores potential extensions for future research within this domain. A list of common acronyms used in this article is presented in Table I.
II IoT Technologies for Wildfire Sensing
@@ II-A1 Visible Light Range Cameras
Numerous research studies have integrated visible range (RGB) camera systems into ground IoT sensing systems to enable fast and robust wildfire detection. Unlike long-range and high-resolution cameras used on watchtowers or satellites, IoT ground sensing systems require small, cost-effective and energy-efficient cameras for scalable deployment. Recent advances in semiconductor technologies and compact lens systems have made it feasible to miniaturise camera modules. To make it more feasible for IoT applications, these low-performance camera modules have also been modified to operate at smaller viewing angles, lower resolutions and frame rates, to reduce energy consumption and the need for large lens systems. For example, these camera modules only provide video quality at a maximum of 1080p30fps with 102 viewing angle. Such camera modules have been used in various IoT projects, including those involving wildfire detection [38, 39, 40, 41, 42, 43, 44]. Notable examples of IoT cameras include IMX219 [45] and OV7670 [46]. Using these IoT cameras, microcontrollers can capture 2D video streams or images in pixel-matrix format, enabling vision-based fire detection. However, it is important to recognise that camera images are sensitive to adverse weather conditions and obstructions and are less reliable in dim conditions, particularly for early fire detection. Additionally, the limited viewing angle resulting from camera miniaturisation requires the use of a rotating platform or multiple cameras to achieve a 360 view if the detection system is based solely on cameras.
@@ II-A2 Infrared Range Cameras
To improve the night visibility of these ground-sensing devices, the implementation of infrared (IR) cameras has been explored as an alternative solution. This approach of paralleling the utilisation of IR technology with visible light cameras has shown a significant improvement in night visibility in watchtowers and has been adopted by many commercial products in fire detection [31, 47]. IR cameras capture thermal images, where each pixel directly represents the temperature or heat intensity of the corresponding area. This characteristic simplifies the fire detection process, providing an intuitive indication of the presence of a fire. In an extremely simplified scenario, fire detection can be achieved by analysing the histogram or sum of the pixel intensities. However, the adoption of IR cameras also comes with a drawback, as they tend to be more expensive compared to regular cameras due to their limited market demand. Although IR cameras facilitate fire detection at night, their reliability can be compromised if the detection distance and temperature range of IR sensors are insufficient. For example, certain IoT compatible IR cameras may only detect temperature ranges from 0-80 C, while forest soil can reach 70 C during high-hazard weather conditions. Therefore, relying solely on IR cameras in an IoT system may not be the optimal choice. In [48], authors proposed a fire detection system that can be applied to IoT or UAVs, and the experimental results show its feasibility by combining the IR camera and the DL detection algorithm.

### v2

# A Survey on IoT Ground Sensing Systems for Early Wildfire Detection: Technologies, Challenges and Opportunities

@@ Abstract
Wildfire, also widely known as bushfire in Australia and forest fire in Europe, present a significant risk to people, the environment and assets. As of 2023, there has been an increase of C in the global surface air temperature compared to the 1850-1900 baseline, attributed to climate change, leading to more days with high fire risk, additionally, the ongoing El Nio event has exacerbated conditions, elevating temperatures to 1.6∘C above pre-industrial levels by early 2024 [1]. As fire weather worsens, fire occurrence increases [2, 3] and the window of opportunity for successful containment and suppression after ignition narrows [4, 5]. Hence, rapid detection of new ignitions will become an increasingly important strategy for fire managers as the climate warms.
@@ I-A Impact of Wildfire Across the Globe
During the 2019-2020 period in Australia, widespread fires known as the “Black Summer” affected various states, burning approximately 24 million hectares, resulting in 33 deaths and around 450 injuries due to smoke inhalation [6]. Besides, the impact on wildlife was immense. According to the World Wide Fund for Nature, billions of animals were injured, killed or displaced [7] due to the Black Summer. For the already fragile ecosystem of animals in Australia, this is a major hit, as many animals, including koalas, wombats and wallabies, are listed as endangered or vulnerable [8].
@@
Worldwide, wildfires are intensifying. According to the European Forest Fire Information System, in 2023 alone, wildfires burned an estimated 908,368 acres of land in countries of the European Union, with regions such as Greece experiencing severe damage and around 20 million tonnes of carbon dioxide emitted [9]. In June 2023, major wildfires erupted in the northeastern areas of Kazakhstan, resulting in the highest annual death toll. At least 15 people lost their lives while fighting the fires [10]. In 2020, the United States experienced the most devastating wildfires in recent history, with more than 50,000 fires burning approximately 3.64 million acres of land, which is the highest figure recorded in the 24-year period from 2000 to 2024. [11]. Despite the consistent number of wildfires in the past three decades, the annual burnt area has increased by 1.62 million hectares from the levels seen in the 1990s in the United States [12]. Similarly, Canada experienced a significant number of wildfires, with around 5,500 fires destroying 17.34 hectares of land [13]. The Brazilian National Institute of Space Research has revealed 3.36 million hectares of land burnt in Brazil annually since 2003 [14]. In addition to these observed increases in fire activity, predicted warming and drying of the climate across much of the world are likely to increase fire activity and result in long-term declines in forest carbon storage, further warming the climate [15]. In light of the increasing threat posed by wildfires, the imperative to develop precise and reliable detection technologies has taken precedence.
Existing Works | Summary | Platforms | ||||
Satellite Remote Sensing | UAV | Camera Watchtower | IoT | |||
Panagiotis et. al. [16] | A review and comparison of various imaging-based technologies for early detection of wildfires | ✓ | ✓ | ✓ | ✓ | |
Ankita et. al. [17] | A review of different technologies for early wildfire detection | ✓ | ✓ | ✓ | ✓ | |
and the comparison of various commercially available early wildfire detection systems | ||||||
Rafik et. al. [18] | An overview of different imaging-based technologies used for the early detection of wildfires | ✓ | ✓ | ✓ | ||
and the application of DL algorithms in fire imaging detection | ||||||
@@
Ahmad et. al. [19] | A review and comparison of various technologies used for early wildfire detection | ✓ | ✓ | ✓ | ||
Faroudja et. al. [20] | A review and comparison focusing on ML/DL algorithms suitable for early detection of wildfires | |||||
Francesco et. al. [21] | A review of ML/DL algorithms suitable for imaging-based early detection of wildfires | ✓ | ✓ | |||
Abdelmalek et. al. [22] | Comparison of ML/DL algorithms suitable for early wildfire detection on UAVs | ✓ | ||||
Hong et. al. [23] | Comparison of ML/DL algorithms suitable for early wildfire detection on camera watchtowers | ✓ | ||||
Mounir et. al. [24] | A review on early wildfire detection using IoT systems | ✓ | ||||
and discussion on challenges in IoT wildfire detection | ||||||
This article | An overview of IoT sensing technologies suitable for early wildfire detection | ✓ | ||||
and the analysis of the challenges in implementing such a system |
@@
I-B Early Wildfire Detection: Existing approaches
The ability to detect wildfires in their early stages allows effective and efficient fire suppression [25]. Specifically, early detection of initial fire stages with benign fire behaviour, enables the timely deployment of fire suppression tactics to increase the probability of initial attack success [5]. Therefore, it is vital to develop early wildfire detection methods and technologies. Here, we present a set of requirements to gauge the capability of existing technologies for early wildfire detection. The algorithm should be able to identify fires while they are still small enough to be successfully contained or suppressed. As the fire grows larger, it may become too difficult to extinguish [5]. Therefore, early detection that can reduce the time it takes to detect and locate small fires is hugely beneficial. Moreover, wildfire detection systems based on environmental monitoring are highly desirable due to the correlation between fire hazards and changes in the environment. Finally, the scalability of wildfire detection systems is beneficial so that systems can not only be deployed on a large scale but also be deployed in different terrains.
We hereby briefly introduce the existing technologies for wildfire detection and analyse their suitability for early wildfire detection. For nearly a century, watchtowers and public reporting have been used to detect wildfires. Unfortunately, watchtowers require observers to remain on top of the tower during certain hours of the day and specific periods of the year, constantly looking for smoke that implies fire and the effectiveness of public reporting is significantly influenced by the population density of the region. With the advancement of computer vision technologies [26, 27, 28], fire detection from watchtowers can now be accomplished day and night with the help of arrays of high-resolution vision / thermal cameras that record visual information, which is then processed using different detection algorithms to recognise smoke, flames, or heat from fire. However, there are some challenges associated with this technology for early wildfire detection. Firstly, fires take some time before producing a signal that is strong enough to detect from a distance, which can prolong their detection time. Secondly, monitoring hilly terrain becomes difficult due to line-of-sight blockage. Consequently, it may be difficult to detect small fires in these areas. Thirdly, this technology lacks the ability to convey environmental changes, such as changes in relative humidity and temperature, which cause diurnal changes in fuel moisture and associated fire activity [29].
As another detection technology, the deployment of crewed fire-spotting aircraft over lightning strike areas or areas of high fire risk is also used, but is often delayed due to resourcing issues or restrictions on night-flying operations. The use of long-endurance Unmanned aerial vehicles (UAVs) equipped with thermal cameras is on the rise, mainly due to their availability to operate in difficult weather conditions, day and night, helping to overcome some of the limitations associated with traditional fire-spotting aircraft [30, 31, 32]. UAVs offer the potential to rapidly confirm suspected ignitions and provide information on the size and behaviour of fires, particularly fires in remote locations. However, authorisation is required and regulations must be followed to operate this system to detect potential fires.
For many years, satellite imagery has been a valuable tool for monitoring large-scale wildfires mainly based on thermal anomalies detected by thermal sensors [33]. But it also comes with limitations, particularly in terms of temporal or spatial resolution of such sensor systems, making it challenging to detect fires when they are still small. The Low Earth Orbit satellite constellation network offers relatively high spatial resolution capable of detecting small fires [34]. However, the temporal frequency is entirely dependent on the size of the constellation, and ensuring prompt detection can be expensive due to the costs associated with deploying and maintaining a large constellation. In addition, opting for a smaller constellation may compromise the reliability of fire detection [35]. Novel algorithms developed specifically for early fire detection using geostationary satellites, which offer temporal resolution of minutes and spatial resolution of kilometres, may offer advances, but have not yet been thoroughly tested [36].
In addition to the technologies mentioned above, a promising complementary approach to detect early wildfires is to use ground sensing systems based on the Internet of Things (IoT). These systems provide real-time data from a variety of ground sensors at desired densities for timely and accurate fire detection, due to the proximity of the ground sensors to the fire ignition locations. This IoT-based system can autonomously raise early alarms for wildfire detection, identify its location, and provide situational awareness by reporting the fire movement. The communication delay from data generation at the ground sensors to reception at the central node can be in a few minutes. Moreover, this kind of system can complement other fire detection and extinguishing systems by identifying the area of interest, which then can be investigated for further details and validation of ignitions as well as to initialise responses, or to monitor areas that are difficult to monitor with other technologies. The benefits of a ground sensing systems would also go beyond wildfire detection and suppression, as they can also be used for environmental monitoring purposes such as fuel moisture monitoring. Due to these unique advantages, many IoT companies have integrated IoT-based wildfire detection systems into their offerings [37, 38, 39, 40]. For example, N5 Sensors offers a system that uses their cutomised gas sensors to identify various volatile compounds, carbon monoxide, and CO2, with the commitment to detect fire ignitions in five minutes [40]. Another example is presented by Dryad, where the Bosch BME680 sensor is used in its IoT setup to detect wildfires, ensuring detection times of under an hour [37]. However, it is essential not to rely solely on one detection system for early fire detection, but to recognise the benefits of combining different wildfire detection approaches and take a more comprehensive approach. This requires long-term research and development efforts across academia and industry on a large scale. In Australia, a leading effort on this front is the recent establishment of the Bushfire Research Centre of Excellence supported by the Australian National University (ANU) and Singtel Optus Pty. Ltd. The centre aims to develop an integrated and layered solution to detect and extinguish small fires using multiple technologies, some mentioned above [25, 32].
I-C Prior Work and Our Contributions
Wildfire detection has been an active topic of research over the past two decades, and several surveys have been published on this general topic. These are summarised in Table I, which highlights the objectives of each survey paper and the specific technological areas on which they focus. Many survey papers, such as [16, 17] and other survey papers cited within these works, provide information on various technologies applicable to early wildfire detection. These surveys evaluate the strengths and weaknesses of different technologies, including the integration of Machine Learning (ML) and Deep Learning (DL) methods for early wildfire detection and prevention. Other survey papers, such as [16, 22, 21] and [23], cover satellite-, UAV- and camera-watchtower-based detection techniques, and few survey papers do cover ground-based IoT systems for early wildfire detection [17, 19, 18, 24, 21, 20]. In [17, 19, 18], summaries of recent advances in early fire detection and prevention technologies are provided, with minimal coverage of IoT systems. In [24], reviews of IoT-based wildfire detection using ML and DL approaches are discussed, together with challenges in IoT wildfire detection. However, in [17, 19, 18, 21, 20] and [24], the challenges faced in the deployment of IoT sensing systems are only briefly mentioned. This is an important gap in the literature that this article aims to address.
In this article, our objective is to explore the potential benefits of modern sensing technologies and their compatibility with IoT sensing systems to detect early-stage wildfires. Specifically, we investigate the use of vision-based and environmental monitoring technologies, combined with anomaly detection algorithms, to detect wildfires in their early stages. Furthermore, the study highlights critical factors and challenges to consider when implementing IoT-based sensing systems for reliable and efficient early wildfire detection, such as sensor placement and network connectivity. The objective of this article is to provide a comprehensive understanding and guidance for the implementation of IoT ground sensing systems for early wildfire detection, together with its challenges and possible research extensions.
The contributions of this work are summarised as follows:
An overview of existing detection techniques using vision and environmental sensing data.
Identification of operational challenges associated with the deployment of IoT systems in remote areas.
Recommendations for developing IoT ground sensing systems in early wildfire detection, along with suggestions to help decision makers identify appropriate deployment strategies according to specified performance metrics.
I-D Paper Organisation
The structure of this article is outlined in Figure 1. Section II discusses the architecture of IoT systems and IoT camera modules and environmental monitoring sensors that can be used for such an application. It also provides a brief overview of the operating principles of different types of sensors, their characteristics and use case examples. Finally, the performance of the sensors is compared in multiple aspects. In Section III, we provide an overview of the wildfire detection algorithms commonly used in IoT systems. These algorithms can be divided into two categories: vision-based detection and anomaly detection based on environmental sensing data. We further classify these detection methods based on their complexity, such as feature detection, time series pattern detection and ML & DL detection in vision-based and threshold detection, time series detection, ML & DL detection in environmental sensing. Additionally, we discuss energy consumption, detection delay, requirements for deploying detection algorithms and the popularity of different detection algorithms is also provided at the end of the section. Section IV explains the operational challenges associated with the deployment of IoT applications for wildfire detection, including sensor limitations, energy efficiency considerations, network coverage constraints, the accuracy of detection algorithms and the challenges related to system deployment. Finally, Section V proposes suggestions for the design and deployment of IoT ground sensing systems in the context of early wildfire detection, offering solutions to most of the challenges identified in Section IV. Recognising the inherent trade-offs among some of these challenges, this section also introduces a performance metric that can guide IoT designers in their decision-making processes. Furthermore, it explores potential extensions for future research within this domain. The list of common acronyms used in this article is presented in Table II.
II IoT Technologies for Wildfire Sensing
@@ II-A1 Visible Light Range Cameras
Numerous research studies have integrated visible range (RGB) camera systems into ground IoT sensing systems to enable fast and robust wildfire detection. Unlike long-range and high-resolution cameras used on watchtowers or satellites, IoT ground sensing systems require small, cost-effective and energy-efficient cameras for scalable deployment. Recent advances in semiconductor technologies and compact lens systems have made it feasible to miniaturise camera modules. To make it more feasible for IoT applications, these low-performance camera modules have also been modified to operate at smaller viewing angles, lower resolutions and frame rates, to reduce energy consumption and the need for large lens systems. For example, these camera modules only provide video quality at a maximum of 1080p30fps with 102∘ viewing angle. Such camera modules have been used in various IoT projects, including those involving wildfire detection [42, 43, 44, 45]. Notable examples of IoT cameras include IMX219 [46] and OV7670 [47]. Using these IoT cameras, microcontrollers can capture 2D video streams or images in pixel-matrix format, enabling vision-based fire detection. However, camera images are sensitive to adverse weather conditions and obstructions and are less reliable in dim conditions, particularly for early fire detection. Additionally, the limited viewing angle resulting from camera miniaturisation requires the use of a rotating platform or multiple cameras to achieve a 360∘ view if the detection system is based solely on cameras.
@@ II-A2 Infrared Range Cameras
To improve the night visibility of these ground-sensing devices, the implementation of infrared (IR) cameras has been explored as an alternative solution. This approach of paralleling the utilisation of IR technology with visible light cameras has shown a significant improvement in night visibility in watchtowers and has been adopted by many commercial products in fire detection [39, 48]. IR cameras capture thermal images, where each pixel directly represents the temperature or heat intensity of the corresponding area. This characteristic simplifies the fire detection process, providing an intuitive indication of the presence of a fire. In an extremely simplified scenario, fire detection can be achieved by analysing the histogram or sum of the pixel intensities. However, the adoption of IR cameras also comes with a drawback, as they tend to be more expensive compared to regular cameras due to their limited market demand. Although IR cameras facilitate fire detection at night, their reliability can be compromised if the detection distance and temperature range of IR sensors are insufficient. For example, certain IoT compatible IR cameras may only detect temperature ranges from 0-80 ∘C, while forest soil can reach 70 ∘C during high-hazard weather conditions. Therefore, relying solely on IR cameras in an IoT system may not be the optimal choice. In [49], authors proposed a fire detection system that can be applied to IoT or UAVs, and the experimental results show its feasibility by combining the IR camera and the DL detection algorithm.